# DOS: Directional Object Separation in Text Embeddings for Multi-Object Image Generation

### v2

@@ C.1 Combining DOS with Attend-and-Excite
Since DOS only modifies CLIP text embeddings, it can be used in combination with other latent modification methods such as Attend-and-Excite (A&E) (Chefer et al. 2023). This combined method addresses some issues that cannot be solved by using either DOS or A&E alone, as shown in Figure 9. We further validate the effectiveness of this combination through a quantitative comparison, presented in Table 10.
Figure 9: Our method, DOS, can be used in combination with other latent modification method such as A&E (Chefer et al. 2023). All methods are based on SDXL.
@@ C.6 Validation with multiple evaluators
While the Success Rate (SR) and Mixture Rate (MR) in Table 1 were measured using GPT-4o-mini (OpenAI 2024), relying on a single Vision Language Model (VLM) could introduce evaluation biases. To validate the reliability of our evaluations, we re-evaluated all methods using two additional evaluators: another VLM, Gemini-2.0-Flash-Lite (Google 2025), and an open-vocabulary object detector, Grounding-DINO (Liu et al. 2024). The MR metric was excluded for the object detector as it is not capable of assessing object mixing. The re-evaluated quantitative comparison, presented in Table 13, shows a consistent trend with our main results, suggesting their validity.
@@
Our main comparison in Table 1 was conducted on SDXL as the base model. To verify that this performance advantage extends to more recent architectures, we compared our method against TEBOpt (Chen et al. 2024) and Self-Cross (Qiu, Wang, and Tang 2025) using SD3.5 as the base model. Table 14 shows that our approach maintains its superior performance on this newer architecture.
@@ C.8 Additional qualitative comparisons
Although our main benchmarks focus on failures arising from inter-object relationships in attribute‑free prompts, we additionally demonstrate that DOS maintains and can even improve attribute binding on the widely used Attend-and-Excite (A&E) benchmark (Chefer et al. 2023). The benchmark defines three prompt templates: (i) “a {animalAanimal~A} and a {animalBanimal~B}”, (ii) “a {animalanimal} and a {colorcolor} {objectobject}”, and (iii) “a {colorAcolor~A} {objectAobject~A} and a {colorBcolor~B} {objectBobject~B}”, of which template (iii) is the most challenging because the model must correctly bind two independent color–object pairs in a single image. Figures 12 and 13 compare DOS with four baselines on template (iii): SDXL, TEBOpt, A&E, and CONFORM. Despite the absence of an explicit attribute–object binding guidance, DOS consistently keeps each color attached to the correct object and delivers clearer separation between the two objects, whereas competing methods often fail to bind attributes correctly or neglect the objects. Although the separation vectors in DOS were devised to mitigate inter‑object interference, they also implicitly preserve the binding between attributes and objects. These results suggest that DOS can be applied effectively to general prompts containing multiple attribute-object tuples, extending its utility beyond specific scenarios.
@@ Appendix D Directional Encoding Across CLIP Text Embedding Types
One of key motivations for our method is the observation that the difference between two CLIP text embeddings encodes directional information, such as $king-man+woman\approx queen$” (Hu et al. 2024a). However, prior work has only validated this directional encoding when semantic token and EOT embeddings are used together. Since DOS calculates separation vectors independently for different embedding types—the semantic token for each object noun and the EOT/pooled embedding—here we test whether this directional property holds for each type of embedding individually.
@@ Appendix E Training Data Biases Underlying Generation Failures
The representation of the trained models are known to be significantly influenced by the distribution of their training dataset (Nguyen et al. 2022; Qin et al. 2023; Xu et al. 2024; Khalafi, Ding, and Ribeiro 2024). Frequent failures in the two problematic scenarios, Dissimilar Background Biases and Many Objects, can be partially explained by this effect. To investigate this, we analyze all captions in LAION 400M (Schuhmann et al. 2021), a dataset whose filtered subset is believed to be used in training T2I models. Specifically, we compare (i) the co-occurrence rate of object pairs with similar vs. dissimilar background biases, and (ii) the proportion of captions containing varying numbers of noun chunks.
@@
In Table 15, we report the co-occurrence rates for two sets of object pairs: one with similar background biases, and the other with dissimilar biases. The co-occurrence rate is defined using the Jaccard index, calculated as the number of captions containing both objects divided by the number of captions containing at least one of them. The results show that object pairs with dissimilar background biases co-occur significantly less frequently than those with similar biases. In Table 16, we examine how the number of noun chunks varies across prompts. Using spaCy (Honnibal et al. 2020) to count noun chunks per caption, we find that over 63% of captions contain fewer than three noun chunks. Together, these findings provide evidence that the failures of T2I models in these two scenarios are, at least in part, attributable to biases in the training data distribution.
Table 16: Distribution of noun chunk counts in LAION 400M captions. The proportion of captions with varying numbers of noun chunks, extracted using spaCy (Honnibal et al. 2020), shows that over 63% of captions contain fewer than three noun chunks, suggesting limited multi-object descriptions in the training data.

### v3

@@ C.1 Combining DOS with Attend-and-Excite
Since DOS only modifies CLIP text embeddings, it can be used in combination with other latent modification methods such as Attend-and-Excite (A&E) (chefer2023attend). This combined method addresses some issues that cannot be solved by using either DOS or A&E alone, as shown in Figure 9. We further validate the effectiveness of this combination through a quantitative comparison, presented in Table 10.
Figure 9: Our method, DOS, can be used in combination with other latent modification method such as A&E (chefer2023attend). All methods are based on SDXL.
@@ C.6 Validation with multiple evaluators
While the Success Rate (SR) and Mixture Rate (MR) in Table 1 were measured using GPT-4o-mini (openai_gpt-4o-mini_2024), relying on a single Vision Language Model (VLM) could introduce evaluation biases. To validate the reliability of our evaluations, we re-evaluated all methods using two additional evaluators: another VLM, Gemini-2.0-Flash-Lite (google_gemini-2.0-flash-lite_2025), and an open-vocabulary object detector, Grounding-DINO (liu2024grounding). The MR metric was excluded for the object detector as it is not capable of assessing object mixing. The re-evaluated quantitative comparison, presented in Table 13, shows a consistent trend with our main results, suggesting their validity.
@@
Our main comparison in Table 1 was conducted on SDXL as the base model. To verify that this performance advantage extends to more recent architectures, we compared our method against TEBOpt (chen2024cat) and Self-Cross (qiu2025self) using SD3.5 as the base model. Table 14 shows that our approach maintains its superior performance on this newer architecture.
@@ C.8 Additional qualitative comparisons
Although our main benchmarks focus on failures arising from inter-object relationships in attribute‑free prompts, we additionally demonstrate that DOS maintains and can even improve attribute binding on the widely used Attend-and-Excite (A&E) benchmark (chefer2023attend). The benchmark defines three prompt templates: (i) “a {animalAanimal~A} and a {animalBanimal~B}”, (ii) “a {animalanimal} and a {colorcolor} {objectobject}”, and (iii) “a {colorAcolor~A} {objectAobject~A} and a {colorBcolor~B} {objectBobject~B}”, of which template (iii) is the most challenging because the model must correctly bind two independent color–object pairs in a single image. Figures 12 and 13 compare DOS with four baselines on template (iii): SDXL, TEBOpt, A&E, and CONFORM. Despite the absence of an explicit attribute–object binding guidance, DOS consistently keeps each color attached to the correct object and delivers clearer separation between the two objects, whereas competing methods often fail to bind attributes correctly or neglect the objects. Although the separation vectors in DOS were devised to mitigate inter‑object interference, they also implicitly preserve the binding between attributes and objects. These results suggest that DOS can be applied effectively to general prompts containing multiple attribute-object tuples, extending its utility beyond specific scenarios.
@@ Appendix D Directional Encoding Across CLIP Text Embedding Types
One of key motivations for our method is the observation that the difference between two CLIP text embeddings encodes directional information, such as $king-man+woman\approx queen$” (hu2024token). However, prior work has only validated this directional encoding when semantic token and EOT embeddings are used together. Since DOS calculates separation vectors independently for different embedding types—the semantic token for each object noun and the EOT/pooled embedding—here we test whether this directional property holds for each type of embedding individually.
@@ Appendix E Training Data Biases Underlying Generation Failures
The representation of the trained models are known to be significantly influenced by the distribution of their training dataset (nguyen2022quality; qin2023class; xu2024demystifying; khalafi2024constrained). Frequent failures in the two problematic scenarios, Dissimilar Background Biases and Many Objects, can be partially explained by this effect. To investigate this, we analyze all captions in LAION 400M (schuhmann2021laion), a dataset whose filtered subset is believed to be used in training T2I models. Specifically, we compare (i) the co-occurrence rate of object pairs with similar vs. dissimilar background biases, and (ii) the proportion of captions containing varying numbers of noun chunks.
@@
In Table 15, we report the co-occurrence rates for two sets of object pairs: one with similar background biases, and the other with dissimilar biases. The co-occurrence rate is defined using the Jaccard index, calculated as the number of captions containing both objects divided by the number of captions containing at least one of them. The results show that object pairs with dissimilar background biases co-occur significantly less frequently than those with similar biases. In Table 16, we examine how the number of noun chunks varies across prompts. Using spaCy (honnibal2020spacy) to count noun chunks per caption, we find that over 63% of captions contain fewer than three noun chunks. Together, these findings provide evidence that the failures of T2I models in these two scenarios are, at least in part, attributable to biases in the training data distribution.
Table 16: Distribution of noun chunk counts in LAION 400M captions. The proportion of captions with varying numbers of noun chunks, extracted using spaCy (honnibal2020spacy), shows that over 63% of captions contain fewer than three noun chunks, suggesting limited multi-object descriptions in the training data.